OpenStudy (anonymous):

solve the equation -6p - 21=3p - 12

OpenStudy (anonymous):

you want to get the variable p to be positive so you move the 6p to the right and the -12 to the left and you get -21+12=3p+6p simplify and you get -9=9p solve for your answer

OpenStudy (anonymous):

p=-1

OpenStudy (anonymous):

-6p - 21 = 3p - 12 (add 21 to both sides) -6p - 21 + 21 = 3p - 12 + 21 (combine like terms) -6p = 3p + 9 (subtract 3p from both sides) -6p - 3p = 3p - 3p + 9 (combine like terms) -9p = 9 (divide by -9 to get p by itself) -9p/-9 = 9/-9 p = - 1 you can now check your answer by subbing in -1 for p -6p - 21 = 3p - 12 -6(-1) - 21 = 3(-1) - 12 6 - 21 = -3 - 12 - 15 = - 15 (correct) ANSWER : p = - 1